### v2

# Interpreting the Linear Structure of Vision-language Model Embedding Spaces

Abstract
Vision-language models encode images and text in a joint space, minimizing the distance between corresponding image and text pairs. How are language and images organized in this joint space, and how do the models encode meaning and modality? To investigate this, we train and release sparse autoencoders (SAEs) on the embedding spaces of four vision-language models (CLIP, SigLIP, SigLIP2, and AIMv2). SAEs approximate model embeddings as sparse linear combinations of learned directions, or “concepts”. We find that, compared to other methods of linear feature learning, SAEs are better at reconstructing the real embeddings, while also able to retain the most sparsity. Retraining SAEs with different seeds or different data diet leads to two findings: the rare, specific concepts captured by the SAEs are liable to change drastically, but we also show that the key commonly-activating concepts extracted by SAEs are remarkably stable across runs. Interestingly, while most concepts are strongly unimodal in activation, we find they are not merely encoding modality per se. Many lie close to – but not entirely within – the subspace defining modality, suggesting that they encode cross-modal semantics despite their unimodal usage. To quantify this bridging behavior, we introduce the Bridge Score, a metric that identifies concept pairs which are both co-activated across aligned image-text inputs and geometrically aligned in the shared space. This reveals that even unimodal concepts can collaborate to support cross-modal integration. We release interactive demos of the SAEs for all models, allowing researchers to explore the organization of the concept spaces.
Overall, our findings uncover a sparse linear structure within VLM embedding spaces that is shaped by modality, yet stitched together through latent bridges—offering new insight into how multimodal meaning is constructed.
footnotetext: These authors contributed equally to this work.
1 Introduction
How do vision-language models (VLMs) organize their internal space in order to relate text and image inputs? Multimodal models encode images and text in a joint space, enabling many impressive downstream multimodal applications. In this paper, we explore how multimodal models encode meaning and modality in order to try and understand how cross-modal meaning is expressed in the embedding space of vision-language models.
We approach this question through dictionary learning: the class of methods that consists of finding linear directions (or, “concepts”) in the latent space of the model, that can break down each embedding into a linear combination of more interpretable directions. We train multiple dictionary learning methods on top of the four VLMs under consideration (CLIP (Radford et al., 2021), SigLIP (Zhai et al., 2023) SigLIP2 (Tschannen et al., 2025) and AIMv2 (El-Nouby et al., 2024; Fini et al., 2024)), and settle on BatchTopK Sparse Autoencoders (SAEs) as the method that dominates the frontier of the tradeoff between remaining faithful to the original embedding space while being maximally sparse. Using the concepts extracted by BatchTopK SAEs, we run a series of investigations to understand how vision-language embedding spaces work. Our analysis is grounded in an ambitious empirical effort: we train and compare SAEs across four distinct vision-language models, using hundreds of thousands of activations, and carry out a thorough suite of evaluations that dissect the geometric, statistical, and modality-related structure of their embedding spaces.
  • We develop VLM-Explore (https://vlm-concept-visualization.com/), an interactive visualization tool that lets researchers explore the linear concepts in a model. VLM-Explore helps visualize which concepts are important in achieving the model’s cross-modal spatial organization by visualizing our BridgeScore metric between pairs of concepts, quantifying how much they contribute to making a joint vision-language space (Section 4)
  • We show that SAEs trained with different seeds and with different training data mixtures are stable and robust when we consider the concepts that are commonly used (high-energy concepts), but very unstable on the concepts that are used very rarely (low-energy concepts) (Section 5).
  • We show that almost all concepts are single-modality concepts: they activate almost exclusively either for text or for image concepts, and this holds across all models that we try (Section 6)
  • However, we also show that this does not mean that the concepts lie within the separate cones of the image and text activations: across models, a large proportion of concepts are almost orthogonal (but not totally orthogonal) to the subspace that encodes modality, and therefore encode largely cross-modal dimensions of meaning (Section 6).
Overall, our findings show how the linear structure of vision-language embedding space can be used to understand the mechanics of joint vision-language space, and we provide the SAEs and the visualization tools for researchers to further explore these connections. Our exploration into the workings of VLM spaces reveals an interesting two-sided conclusion: while the space is organized primarily in terms of modality, single-modality can still be related through high cosine similarities on the subspace that is orthogonal to modality, creating cross-modal bridges of meaning.
2 Background and related work
Linear Representations in Neural Networks
The representation spaces of neural networks have been often shown to encode many important features (like syntactic or object-category information) largely linearly (Alain & Bengio, 2016; Hewitt & Manning, 2019; Belinkov, 2022; Li et al., 2016; Conneau et al., 2018; Fel et al., 2024a), since, after all, the operations on these spaces are largely linear In vision-language models, strong performance on zero-shot tasks suggests that their embeddings are structured along meaningful directions. Empirical evidence from probing and representation analyses further supports the idea that semantic content is linearly organized Muttenthaler et al. (2023; 2024); Nanda et al. (2021); Tamkin et al. (2023); Bricken et al. (2023).
Sparse Coding and Dictionary Learning
Dictionary Learning (Tovsić & Frossard, 2011; Rubinstein et al., 2010; Elad, 2010; Mairal et al., 2014; Dumitrescu & Irofti, 2018) has emerged as a foundational framework in signal processing and machine learning for uncovering latent structure in high-dimensional data. It builds upon early theories of Sparse Coding, developed in computational neuroscience Olshausen & Field (1996; 1997); Foldiak & Endres (2008), where sensory inputs were modeled as sparse superpositions of overcomplete basis functions. The core objective is to recover representations where each data point is approximated as a linear combination of a small number of dictionary atoms Hurley & Rickard (2009); Eamaz et al. (2022). This sparsity constraint encourages interpretability by associating individual basis vectors with meaningful latent components. In neural network interpretability, sparse decompositions have gained traction as tools to extract interesting features from learned representations (Elhage et al., 2022; Cunningham et al., 2023; Fel et al., 2023; 2024b; Rajamanoharan et al., 2024; Gorton, 2024; Surkov et al., 2024). Sparse autoencoders (SAEs) (Makhzani & Frey, 2014)  are a scalable instantiation of Sparse Dictionary Learning implemented by training a small neural network. Recent work has shown that SAEs extract semantically meaningful concepts (Bricken et al., 2023; Bussmann et al., 2024; Thasarathan et al., 2025; Paulo & Belrose, 2025; Gao et al., 2025; Bhalla et al., 2024; Fel et al., 2025), often yielding more interpretable and compositional decompositions than PCA or ICA (Bussmann et al., 2024; Braun et al., 2024; Chen et al., 2024; Fel et al., 2024b; Makelov et al., 2023).
Modality and Interpretability in Multimodal Models
A growing body of work has examined how multimodal models encode meaning and modality in their internal representations. Despite being trained to align modalities, vision-language models exhibit a persistent modality gap, where text and image embeddings occupy distinct conical regions in the joint space (Liang et al., 2022; Shukor & Cord, 2024). However, examining the neurons (Goh et al., 2021; Shaham et al., 2024), representations (Bhalla et al., 2024; Parekh et al., 2024; Wu et al., 2024) shows that multimodal models showcase aspects of cross-modal internal processing.
3 Methods: Sparse Autoencoders on Vision-Language Models
Figure 1: Sparsity–Expressivity trade-off. Pareto fronts for five dictionary learning methods applied to four vision-language models (CLIP, SigLIP, SigLIP 2, AIMv2). Each curve shows the trade-off between reconstruction quality ($R^{2}$ score) and sparsity level ($\ell_{0}$ norm of $\mathbf{Z}$). The three SAEs (TopK, JumpReLU, and BatchTopK) consistently achieve the best balance, with BatchTopK slightly dominating other sparse autoencoder variants.
Background and Notation.
We denote vectors by lowercase bold letters (e.g., $\bm{x}$) and matrices by uppercase bold letters (e.g., $\bm{X}$). We write $[n]$ to denote the index set $\{1,\dots,n\}$. The unit $\ell_{2}$-ball in $\mathbb{R}^{d}$ is defined as =$\mathcal{B}=\{\bm{x}\in\mathbb{R}^{d}\mid\|\bm{x}\|_{2}\leq 1\}$. We consider a general multimodal representation learning setting in which a vision-language model maps image or text inputs $\bm{x}\in\mathcal{X}$ into a shared representation space $\mathcal{A}\subseteq\mathbb{R}^{d}$. A collection of $n$ such embeddings is stored in a matrix $\bm{A}\in\mathbb{R}^{n\times d}$. Dictionary Learning seeks to approximate $\bm{A}$ as a linear combination of concept vectors from a dictionary $\bm{D}\in\mathbb{R}^{c\times d}$, using a sparse code matrix $\bm{Z}\in\mathbb{R}^{n\times c}$:
$(\bm{Z}^{\star},\bm{D}^{\star})=\arg\min_{\bm{Z},\bm{D}}\|\bm{A}-\bm{Z}\bm{D}\|_{F}^{2}$ (1)
where $\|\cdot\|_{F}$ denotes the Frobenius norm, and dictionary atoms are constrained to lie on $\mathcal{B}$.
Sparse Autoencoders.
In the case of Sparse Autoencoders (SAEs), an encoder network $\bm{\psi}$, a single-layer MLP, maps embeddings $\bm{A}$ to sparse codes $\bm{Z}$ via a linear transformation and a projection (Hindupur et al., 2025) operator $\bm{\Pi}\left\{\cdot\right\}$:
$\bm{Z}=\bm{\psi}(\bm{A})=\bm{\Pi}\left\{\bm{A}\bm{W}+\bm{b}\right\}$ (2)
where $\bm{W}\in\mathbb{R}^{d\times c}$ and $\bm{b}\in\mathbb{R}^{c}$ are learnable encoder parameters. In particular, we consider several SAEs that can be described by their projection operators such as:
$\bm{\Pi}\left\{\bm{x}\right\}=\begin{cases}\mathrm{ReLU}(\bm{x})=\max(\bm{0},\bm{x})\\[3.00003pt] \mathrm{JumpReLU}(\bm{x})=\max(\bm{0},\bm{x}-\bm{\theta})+\bm{\theta}\odot H(\bm{x}-\bm{\theta})\\[3.00003pt] \mathrm{TopK}(\bm{x},k)=\arg\min_{\bm{z}\in\mathbb{R}^{c}}\|\bm{z}-\bm{x}\|_{2}^{2}\quad\text{s.t.}\quad\|\bm{z}\|_{0}=k,\;\bm{z}\geq\bm{0}\end{cases}$ (3)
Where $H$ is the Heaviside step function. We will also explore a simple extension, BatchTopK, which flexibly shares applies the TopK operation across the entire batch.
Expressivity vs Sparsity.
An ideal dictionary learning method should produce representations that are both expressive and sparse. In practice, this means that model embeddings should activate only a small number of meaningful concepts, indicating that the learned dictionary is both specific and semantically aligned with the data. We apply five dictionary learning methods to four models (CLIP, SigLIP and SigLIP2, which are encoder models, and AIMv2 which is an autregressive vision-language model). The activation matrix $\bm{A}$ consists of 600,000 normalized embeddings from passing the COCO dataset (Lin et al., 2014) through the models, with each $\bm{A}_{i}\in\mathcal{B}$, the unit $\ell_{2}$-ball.
We evaluate all SAEs on their trade-off between sparsity – measured as the $\ell_{0}$ norm of $\bm{Z}$, i.e., the number of non-zero entries per code vector – and reconstruction quality, quantified by the $R^{2}$ score between the original activations $\bm{A}$ and their reconstructions $\bm{Z}\bm{D}$. This trade-off captures how well the learned codes retain information while encouraging interpretability through sparsity. Results, shown in Figure 1, indicate that all four SAEs variants outperform Semi-NMF at fixed sparsity levels, with BatchTopK achieving slightly better performance overall. In practice, we also find that training TopK and BatchTopK SAEs is more stable and requires less hyperparameter tuning. For these reasons, we use the BatchTopK SAE as our default model in all subsequent experiments.
3.1 Metrics: The Geometry and Statistics of Learned Concepts
To assess the interpretability and quality of the learned representations, we introduce four complementary metrics that go beyond the standard sparsity–reconstruction trade-off: stability, energy, modality score, and bridge score. Each of these measures captures a different axis of representational information to better understand the representation of VLM using the SAEs.
Stability.
Stability quantifies the consistency of learned dictionaries under different training runs with random initialization. If two independently trained SAEs on the same model produce significantly different dictionaries, it suggests that the learned concepts are not robust, which weakens their interpretability. Following Fel et al. (2025) and Spielman et al. (2012), we compute the stability between two dictionaries $\bm{D},\bm{D}^{\prime}\in\mathbb{R}^{c\times d}$ by optimally aligning their rows using the Hungarian algorithm, which solves for a permutation matrix $\bm{P}\in\mathcal{P}(c)$ that maximizes the total similarity between matched concept vectors:
$\text{Stability}(\bm{D},\bm{D}^{\prime})=\max_{\bm{P}\in\mathcal{P}(c)}\frac{1}{c}\operatorname{Tr}(\bm{D}^{\top}\bm{P}\bm{D}^{\prime}).$ (4)
A higher stability score indicates that concept representations are reproducible across training runs, reinforcing their semantic reliability.
Energy.
While stability measures whether concepts are consistently learned, it does not tell us which concepts matter most in practice. To better understand which concepts are actually used by the model, we turn to the energy of each concept. Energy captures the average activation strength of a concept across a dataset, revealing which parts of the dictionary are consistently involved in reconstructions. Formally, for concept $i$, we define its energy as: $\text{Energy}_{i}=\mathbb{E}_{\bm{z}}(z_{i})$. Energy provides a statistical view over the dataset, revealing where attention should be focused for interpreting the learned representation.
Modality Score.
In the context of VLMs, a central question is whether a concept is modality-specific or shared across modalities. To probe this, we introduce the modality score, which quantifies how much a concept contributes to reconstructing image versus text inputs. Let $\iota$ and $\tau$ be the empirical distributions of sparse codes $\bm{z}$ from image ($\iota$) and text ($\tau$) inputs, respectively. We define the modality score of a concept $i$ as the fraction of expected activation energy assigned to image-based inputs:
$\text{ModalityScore}_{i}=\frac{\mathbb{E}_{\bm{z}\sim\iota}(z_{i})}{\mathbb{E}_{\bm{z}\sim\iota}(z_{i})+\mathbb{E}_{\bm{z}\sim\tau}(z_{i})}.$ (5)
A modality score near $1$ indicates that the concept fires only image inputs, while a score near 00 suggests text dominance. A value around $\frac{1}{2}$ indicates a multi-modal concept.
Bridge Score.
While the modality score offers a per-concept perspective, many vision–language tasks rely on coordinated interactions between concepts across modalities. To capture these interactions, we introduce the bridge matrix $\mathbf{B}\in\mathbb{R}^{c\times c}$, which quantifies how pairs of concepts contribute to the alignment of modalities. Let $(\bm{z}_{\iota},\bm{z}_{\tau})\sim\gamma$ be a pair of sparse codes obtained from a matching image–text input pair, drawn from a joint empirical distribution $\gamma$ over aligned data points. Let $\bm{D}\in\mathbb{R}^{c\times d}$ denote the shared dictionary. The bridge matrix is defined as:
$\mathbf{B}=\underbrace{\mathbb{E}_{(\bm{z}_{\iota},\bm{z}_{\tau})\sim\gamma}(\bm{z}_{\iota}^{\top}\bm{z}_{\tau})}_{\text{co-activation}}\odot\underbrace{(\bm{D}\bm{D}^{\top}),}_{\text{alignment}}$ (6)
with $\odot$ being the Hadamard product. The first term captures statistical co-activation between concepts: how frequently concept $i$ in the image code and concept $j$ in the text code are simultaneously active for semantically aligned inputs. The second term reflects the geometric alignment between the corresponding dictionary atoms, computed as the cosine similarity between dictionary atoms The elementwise product combines both aspects—activation and alignment—into a single interpretable structure.
The resulting matrix $\mathbf{B}$ reveals how concepts jointly operate across modalities. A large value at position $\mathbf{B}_{i,j}$ indicates that concepts $i$ and $j$ frequently co-activate and encode similar directions in the shared embedding space. These concept pairs act as semantic bridges, facilitating the alignment of image and text embeddings.
In this sense, the bridge matrix provides insight into why the model succeeds at cross-modal alignment. It highlights not just which concepts are shared, but how they collaborate structurally and statistically to support the model’s semantic integration of vision and language.
4 VLM-Explore: An interactive, functional concept explorer
Our first contribution is the development of VLM-Explore, an interactive visualization tool designed to facilitate the exploration and analysis of concept representations extracted from VLM using SAEs. This platform provides a functional interface for examining the internal structure and organization of learned concepts within the joint vision-language embedding space. The core view of VLM-Explore presents a two-dimensional UMAP projection of the SAE dictionary $\bm{D}$, where each point corresponds to a learned concept vector. The visual encoding incorporates both statistical and semantic information: the size of each point represents the energy of the concept, and the color of each point encodes its modality score (Equation 5). This visualization enables intuitive, high-level assessment of how concepts are distributed with respect to modality and influence within the learned representation.
Upon selecting a concept, the interface displays its top activating samples from the COCO dataset, thereby offering concrete examples of the semantic attributes captured by the concept. A central feature of VLM-Explore is its “co-occurrence” view, which renders edges from the selected concept to the concepts with the highest bridge score. The thickness of each edge reflects the magnitude of this score, thereby emphasizing concept pairs that contribute substantially to cross-modal alignment. This interactive graph-based representation offers insight into the latent relational structure of the concept space, revealing both local interactions and global patterns of semantic organization.
By supporting the identification of dominant, modality-specific, and cross-modal bridging concepts, VLM-Explore serves as a diagnostic and interpretive tool for understanding how VLMs organize semantic content across modalities. Concept pairs with high bridge scores, in particular, can be viewed as structural components that facilitate the alignment objective implicitly optimized by the underlying VLM. We release VLM-Explore in conjunction with all trained SAE dictionaries for the four models analyzed in this work—CLIP, SigLIP, SigLIP2, and AIMv2—offering an open platform for further research into the compositional and linear structure of multimodal embedding spaces.
Figure 2: Stability vs. Energy. (Left) Energy is concentrated on a few concepts, with 512 concepts getting 99% of the energy, and the other 3,500 concepts only sharing 1% of the total coefficient weight. (Center) When training with 5 different random seeds, the stability for all 4,096 concepts is low – 0.16 (Right) When we weight concepts by energy, we can see that the concepts that are used often in reconstruction are actually stable, with stability of 0.92 if we take the top 512 energetic concepts from each run. The instability comes from concepts that are rarely if ever actually used.
5 Analysis 1: Are We Finding Consistent Model Features?
Before analyzing the semantic content of the concepts extracted by our SAEs (Section 6), we first examine whether these concepts are consistent—both across training runs with different random seeds, and across variations in the SAE training dataset. We study stability across: (i) random seeds—to test sensitivity to training stochasticity, and (ii) data modality mixtures—to test dependence on the distribution of input activations.
Results: Stability Across Different Seeds
We train five SAEs with different random seeds on the same SigLIP2 activations, each using $4096$ concepts, and compute the average pairwise stability (Equation 4) between them. As shown in Figure 2 (center), at first glance, overall stability is low: the mean similarity between dictionaries is just $0.16$, suggesting high sensitivity to random initialization. However, this apparent instability is misleading. On the left, see that reconstruction energy is heavily concentrated in a small subset of concepts: the top $512$ concepts account for approximately $99\%$ of the total activation mass. To isolate the functionally relevant features, we recompute stability using only the top-$k$ most energetic concepts from each run. We find that high-energy concepts are remarkably stable, with near-perfect alignment across seeds (Figure 2, right). This reveals a key paradox:
Finding: Without conditioning on energy, SAEs concepts appear unstable; but when viewed in conjunction with their associated energy, a consistent and robust set of model features emerges.
Figure 3: UMAP visualization of the SAE concept spaces under different image–text data mixtures. Color indicates the modality score of each concept. The top row displays all concepts with equal size; the bottom row scales point size by concept energy. While the dominant modality in the training data strongly influences the number of concepts recovered (top row), the most energetic concepts (bottom row) remain relatively stable across mixtures. A similar set of high-energy text and image concepts is recovered regardless of the input distribution. Additional results for other models are provided in Appendix A, Figure 9.
The instability arises almost entirely from the low-energy tail of the dictionary—concepts that contribute little to the reconstruction objective. We conclude that SAEs trained with different seeds reliably recover a core set of functional, high-usage concepts.
Results: Stability Across Different Training Data Mixtures
The analysis above addresses random variation, but it remains unclear whether the recovered concepts reflect structure in the model or simply the statistics of the dataset used to train the SAE. To test this, we vary the modality mixture of the SAE training activations: how many come from image inputs vs. text inputs. We train SAEs on seven data compositions, ranging from heavily image-dominant (5:1 image:text) to heavily text-dominant (1:5). Results are shown in Figure 3. As expected, changing the data mixture shifts the overall modality distribution of the learned concepts. However, when weighting by energy (bottom row), the most energetic concepts remain stable across all mixtures. High-energy directions consistently appear regardless of the training distribution, and exhibit similar activation patterns.
These analyses show that while SAE dictionaries may appear unstable, this instability is concentrated in low-energy concepts. The high-energy subset—responsible for nearly all reconstruction—is both robust to random initialization and is also not simply a reflection of the data. Crucially, this consistency only holds when conditioning on energy, highlighting that stability emerges from how the model uses the learned concepts.
Discussion: Why Low-Energy Concepts?
Our evaluations point to two main findings. First, the important (i.e., high-energy) concepts in our SAEs are stable: they remain consistent across different seeds and are largely unaffected by changes in the image-text data mixture. Second, the low-energy concepts are highly unstable, with different SAEs discovering widely different, rarely-used features. Does this mean that low-energy concepts are useless? Not necessarily. Leveraging VLM-Explore (Section 4) we observe that many low-energy concepts correspond to coherent but rare patterns, such as “yaks” or “valentines day hearts.” These concepts may be semantically meaningful, but their contribution to reconstruction is small, and thus SAEs do not consistently recover them. By contrast, high-energy concepts typically capture more concepts like “red” or “two women” and are consistently recovered because they support a wide range of reconstructions.
These findings guide the rest of this paper. They suggest that SAEs can be used as reliable tools for interpreting VLM embeddings—but only when conditioned on energy. Concepts with high energy form a stable and informative basis; concepts with low energy may still be meaningful, but should be treated with caution due to their instability.
6 Analysis 2: Are We Finding Features With Cross-Modal Meaning?
Having established that SAEs recover stable and consistent high-energy features, we now ask whether these features capture structure that spans both modalities. Specifically, we analyze whether the learned concepts reveal aspects of the geometry of the joint vision-language embedding space, and whether they support cross-modal alignment.
Most Concepts Are Single-Modality
Figure 4: Histogram of the modality scores (top) and the modality scores weighted by energy (bottom) of every concept in the SAEs that we trained for all four models. On every model, the modes are the two extremes: concepts either activated only by text or only by image. However, weighting by energy makes this much more prominent, showcasing that almost all of the reconstruction coefficients go to concepts that are single-modality.
Our first observation is that the majority of concepts activate predominantly on a single modality—either text or image. As defined in Equation 5, the modality score of a concept measures the fraction of its total activation energy that comes from image inputs. A score near 1 indicates image-specific usage; a score near 0 indicates text-specific usage. In Figure 4, we plot the modality score distribution across all concepts for SAEs trained on each model. The top row shows raw counts: while some concepts cluster around 0.5, the distributions are clearly bimodal, with dominant modes at 0 and 1. This pattern becomes more pronounced when we weight each concept by its energy (right): almost all of the energy across all models is concentrated on concepts that are nearly exclusive to one modality.
We conclude that the most influential directions in the SAE dictionary — those responsible for the bulk of reconstruction — are overwhelmingly single-modality. Despite the shared embedding space, the sparse structure learned by the SAE reveals a strong separation between text and image representations.
Cross-Modal Alignment from Unimodal Bases
Interestingly, the predominance of single-modality concepts in the SAE dictionary does not preclude the model’s ability to align across modalities. This sets up a central tension: if the concepts that dominate reconstruction are modality-specific, how does the model support semantic connections between paired image–text inputs? The resolution lies in the structure of the SAE dictionary itself. While most concepts activate primarily on one modality, they do not exist in isolation. The sparse autoencoder enables bridging via geometrically aligned and co-activated concept pairs—concepts that live in distinct modality clusters but consistently respond to semantically related inputs.
Finding: Despite being functionally unimodal, concepts can form semantically meaningful cross-modal bridges through high co-activation and directional alignment. These bridges support joint understanding across text and image even without shared concepts.
As illustrated in Figure 5, high-bridge-score connections often capture many-to-many relationships—for example, a visual concept like “surfboards” aligns not just to a single caption, but to a diverse cluster of related descriptions. This illustrates how the model achieves flexible semantic grounding across modalities using sparse but structured interconnections.
Figure 5: Bridge score identifies semantically aligned concept pairs across modalities. Even though most concepts are unimodal in activation, many form meaningful connections across modalities via the bridge score. Edges connect concept pairs that are both geometrically aligned and co-activated on paired examples. These links often reflect many-to-many relationships between clusters of related concepts, enabling cross-modal alignment despite sparse unimodal structure.
Many Concepts Are Nearly Orthogonal to the Modality Subspace
Given that most concepts are single-modality in usage, one might ask whether they are also geometrically aligned with modality – i.e., whether they lie along directions that distinguish image from text embeddings. If so, this would suggest that the concepts primarily reflect modality separation, rather than shared or cross-modal meaning. We begin by analyzing the embedding space itself. Prior work has shown that representations in deep models often occupy anisotropic regions – narrow cones within the full space (Ethayarajh, 2019; Mimno & Thompson, 2017). In our setting, we ask whether image and text embeddings occupy separate cones (Tyshchuk et al., 2023): we compute the average cosine distance between pairs of embeddings from the same modality versus different modalities. As shown in Figure 6, embeddings of the same modality (green, blue) are significantly closer than those of different modalities (orange). This suggests that image and text activations reside in distinct conical regions of the embedding space.
Figure 6: Within-modality anisotropy for CLIP and SigLIP. Cosine similarities show that image–image and text–text pairs (blue/green) are closer than image–text pairs (orange), indicating that embeddings occupy separate modality-specific cones.
Do the concept vectors align with the modality structure? If they follow the geometry of the modality cones—separating image from text—they likely encode little cross-modal content. But if many lie near directions orthogonal to the modality subspace, they may instead reflect shared meaning. We test this by using each concept as a linear modality classifier: we project activations onto the concept direction and measure how well it separates image from text. High accuracy implies alignment with modality; near-chance performance suggests the concept is modality-agnostic and may encode cross-modal structure.
Figure 7: Histogram of the accuracies of concepts used as classifiers, with the left side weighting each concept in the histogram by its energy. An accuracy of 0.5 means that the concept is orthogonal to every linear direction that defines modality (and therefore likely encodes some aspect of meaning independent from modality), and an accuracy of 1 means that the concept is aligned with a modality direction. We see that, especially in CLIP and AIMv2, there is a significant proportion of concepts that are almost orthogonal to the modality subspace. When we weight by energy, we see that accuracy-1 concepts are generally very high-energy across models, but that concepts less aligned with modality still receive a significant amount of energy.
Figure 7 presents the distribution of classification accuracies for all concepts (left) and weighted by energy (right). While many concepts—particularly high-energy ones – are highly predictive of modality, a substantial number achieve near-random accuracy. These findings suggest that although most energetic concepts are modality-specific in activation, many are geometrically close to a modality-agnostic subspace. This points to the existence of sparse linear directions that support semantic alignment across modalities while remaining sensitive to modality itself.
Figure 8: The SAE projection effect. Even if a concept direction is nearly orthogonal to modality, the SAE encoder projects activations through $\bm{Z}=\bm{\Pi}\left\{\bm{A}\bm{W}+\bm{b}\right\}$, selecting the top-scoring concepts per input. Because the input distributions differ by modality, the thresholding step (e.g., TopK) may consistently select the concept for one modality but not the other. This creates functionally single-modality concepts, even when their directions lie near the modality-agnostic subspace.
How can concepts be both single-modality and almost orthogonal to modality? The SAE projection effect.
The results reveal an apparent contradiction: many concepts activate almost exclusively for one modality (Figure 4), yet their directions are poor modality classifiers (Figure 7), suggesting they lie near the modality-agnostic subspace. How can a concept be geometrically neutral yet functionally single-modality?
The resolution lies in the SAE projection step: activations are encoded as $\bm{Z}=\bm{\Pi}\left\{\bm{A}\bm{W}+\bm{b}\right\}$, where $\bm{\Pi}\left\{\cdot\right\}$ enforces sparsity (e.g., TopK or ReLU). Even if a concept direction is nearly orthogonal to modality, one modality may consistently produce higher pre-activation scores, causing it to be selected more often. This leads to modality-specific usage driven by modality-skewed thresholding, rather than geometry. We illustrate this effect in Figure 8.
Finding: A concept direction can be almost orthogonal to modality in direction, yet be functionally single-modality in activation, due to the interaction between input distribution and the sparsity projection.
7 Discussion and Conclusion
We study the linear structure of vision-language embedding spaces using SAEs trained on four VLMs. While concept dictionaries vary across seeds and data mixtures, the high-energy subset is consistently recovered and accounts for nearly all reconstruction. Most concepts are unimodal in usage but often lie near the modality-agnostic subspace, suggesting shared structure shaped by SAE thresholding rather than direction. To analyze cross-modal alignment, we introduce the Bridge Score, identifying concept pairs that are both geometrically aligned and co-activated. We release VLM-Explore, an interactive tool to visualize these structures across models.
These insights lay a foundation for building more interpretable and controllable multimodal models, and open the door to future research into structured alignment mechanisms.
Acknowledgments
This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University.
Appendix A Additional Figures and Expanded Discussion
We propose to supplements the main analyses by providing extended visualizations and detailed commentary for all four models—CLIP, SigLIP, SigLIP2, and AIMv2—used throughout the paper. Each subsection corresponds to a specific analysis presented in Section 5 and Section 6 of the main paper.
A.1 UMAP Visualizations Across Data Mixtures
Figure 9: UMAP projections of SAE concept dictionaries across various image-text training mixtures for SigLIP2, SigLIP, and CLIP. Top row: equal dot size. Bottom row: dot size scaled by energy. Related to Figure 3.
Figure 9 shows UMAP projections of the learned SAE concept dictionaries across different training data mixtures. Each subfigure corresponds to a different VLM (SigLIP2, SigLIP, CLIP), with UMAP applied to the learned dictionary atoms. These visualizations complement the main Figure 3, now extended to multiple models.
The top rows present concepts with uniform dot size, providing an unweighted view of the dictionary’s structure. Here, modality is encoded by color reflecting the modality score (as defined in Eq. 5 of the main paper). The bottom rows rescale each concept’s dot size according to its energy score, revealing the relative contribution of each concept to the reconstruction task.
Our key observations include:
  • Across all models, when data mixtures are skewed (e.g., more image than text inputs), the distribution of concepts by modality shifts accordingly. This confirms the sensitivity of low-energy, rarely-used concepts to training data statistics.
  • However, the bottom rows show that energy-dominant (i.e., frequently activated) concepts remain consistently placed in similar areas of the space, regardless of training mixture. These concepts represent a stable semantic core that is robust to dataset composition.
This supports our conclusion from Section 5 that the SAE models recover a stable, high-energy subspace, even under strong shifts in input distributions.
A.2 Modality Score Histograms
Figure 10: Modality score histograms for all models. Top: unweighted. Bottom: weighted by energy. Related to Figure 4.
The histograms in Figure 10 provide an empirical distribution over modality scores across all four models. The top row shows raw counts of modality scores without weighting; the bottom row weights each concept by its energy.
We observe two critical patterns:
  • The modality score distribution is clearly bimodal across all models: most concepts are either strongly image-specific (score near 1) or text-specific (score near 0). This confirms that although the joint embedding space is shared across modalities, its internal structure is dominated by modality-specific axes.
  • When weighted by energy, this bimodality becomes more pronounced: nearly all reconstruction-relevant concepts are unimodal. That is, high-energy concepts are not just modality-skewed : they are overwhelmingly unimodal in practice.
These results extend the findings of Section 6 and illustrate that, while multimodal alignment exists, it emerges from indirect interactions among strongly unimodal building blocks.
A.3 Within-Modality Anisotropy
Figure 11: Within-modality anisotropy: Cosine similarity distributions across image-image, text-text, and image-text embedding pairs. Related to Figure 6.
Figure 11 presents the cosine similarity distributions between pairs of embeddings from the same and different modalities for each model. These distributions reveal that image-image and text-text embedding pairs consistently exhibit higher mutual similarity than image-text pairs, supporting the conical anisotropy hypothesis discussed in the main paper.
This structure—text and image activations occupying separate cones—is consistent across CLIP, SigLIP, SigLIP2, and AIMv2. It confirms that the VLMs, despite being trained for modality alignment, still maintain a strongly separable embedding geometry.
A.4 Concept Classification Accuracy
Figure 12: Histogram of concept-wise modality classification accuracy for all models. Top: unweighted. Bottom: weighted by energy. Related to Figure 7.
Figure 12 shows the results of using each concept as a linear modality classifier—testing how well projection onto a concept vector can separate image vs. text embeddings.
Insights from this analysis:
  • Many concepts (particularly low-energy ones) perform at near-random accuracy, suggesting that they lie close to the modality-agnostic subspace and may encode more abstract or shared meaning.
  • High-energy concepts, however, often act as strong modality classifiers (accuracy near 1), especially in CLIP and AIMv2. Yet, importantly, not all high-energy concepts are modality-aligned indicating that semantically meaningful, shared features may still contribute to reconstruction.
These findings support the nuanced picture presented in Section 6: although usage is often unimodal, some concept vectors are not aligned with the modality separation direction, suggesting a functional dissociation between activation and geometric alignment.
Appendix B Illustrative Examples from the Interactive Demo
We now present examples from the VLM-Explore interface (Section 4), illustrating concrete semantic content of discovered concepts and their inter-modal associations via bridge scores.
Figure 13: Interactive Demo: Example of text concept “banana” and associated visual concepts in CLIP, discovered using the bridge score. Shows clear cross-modal alignment despite being distant in UMAP space (non-trivial associations).
Figure 13 shows the concept cluster for the keyword “banana” as rendered in the VLM-Explore interface. Text instances grouped under this concept activate strongly for semantically aligned visual representations. The bridge score identifies connected visual concepts, which often correspond to similar object-level semantics (e.g., bananas on a table, being held, or cut). This demonstrates that the bridge score successfully identifies semantically aligned relations.
Figure 14: More aligned concept examples (CLIP): Object and scene-level text concepts and their linked visual activations. Shows concept-level semantic granularity.
These examples show how the bridge framework recovers granular semantic regularities across modality, spanning categories like scenes, humans, and objects. Despite functional single-modality usage (due to projection effects), the bridge allow to show the practical alignment.
Figure 15: Relational concepts: “cat sitting,” “man skateboarding,” and “holding umbrella.” These examples reflect the model’s ability to capture spatial or action-based semantics.
The bridge score also captures compositionality. In Figure 15, relational and action-based meanings are represented—text concepts describing activities map to corresponding image concepts, indicating joint semantic structure beyond raw modality.
Figure 16: Concepts corresponding to texture and color: “black and white” and “wooden.” These illustrate the fine-grained properties encoded by individual SAE features.
Finally, Figure 16 demonstrates the interpretability of concepts capturing non-object attributes – e.g., texture, material, and tone. These show that the SAE not only captures objects or scenes, but also abstract perceptual features.